OpenStudy (vortish):

solve by any method 4x^2+36X=88

OpenStudy (tkhunny):

First thing I would do is divide everything by 4.

OpenStudy (vortish):

I Know that you get 4(x^2+36x-22)=0 this is the step were that factoring comes in and im geting confused

OpenStudy (campbell_st):

let it equal zero then take a common factor or 4 \[4(x^2 + 9x - 22) = 0\] find the factors of -22 that add to 9 then the solution is 4(x + factor 1)(x + factor 2) = 0 then solve x + factor 1 = 0 x + factor 2 = 0 to get the solutions.

OpenStudy (tkhunny):

I said divide, not factor. Get rid of it. 4 is not ever zero. You're done with it.

OpenStudy (vortish):

what i come up with is X=2 and x = -11

OpenStudy (vortish):

but the 36x is still there and it is devisable by 9

OpenStudy (vortish):

well campbell if you add 2 to -11 you come up with -9

OpenStudy (anonymous):

\[4x^2+36x=88\]subtract the \(88\) and set the equation equal to \(0\):\[4x^2+36x-88=0\]factor out the \(4\):\[4(x^2+9x-22)=0\]the trinomial \(x^2+9x-22\) can be factored out to be:\[4[(x-11)(x+2)]=0\]now set each of the factors equal to \(0\):\[4=0\]\[x-11=0\]\[x+2=0\]and solve for each:\[4\neq0\]\[x=11\]\[x=-2\] tada! :)

OpenStudy (anonymous):

hope that helps! @vortish :)

OpenStudy (campbell_st):

what about 11 + -2 = 9 so you are looking at 4(x - 2)(x + 11) = 0 now solve x - 2 = 0 and x + 11 = 0 for you solutions.